A student wishes to titrate a solution of glycine.HCl with a 0.1000 M solution of NaOH using a 50-mL buret. If the student desires the second equivalence to occur at approximately 30. mL of titrant, what mass of glycine.HCl should be dissolved in water for the titration?
Answer #1

To obtain the second point of equivalence at 30 mL, the first point of equivalence at 15 mL was obtained, the required mass of glycine-HCl is calculated:

m glycine-HCl = M * V * MM = 0.1 M * 0.015 L * 111.53 g / mol = 0.167 g

If you liked the answer, please rate it positively, you would help me a lot, thanks.